The current price of a non-dividend-paying stock is 30. The volatility of the stock is 0.3 per annum. The risk free rate is 0.05 for all maturities. Using the Cox-Ross-Rubinstein binomial tree model with two time steps to do the valuation, what is the value of a European call option with a strike price of 32 that expires in 6 months?
Answer #1

uret dre ovi = 1

u = e^(0.3*(0.25)^(1/2) = 1.1618

d = 1/u = 1/1.1618 = 0.8607

Step 5 Binomial Tree Pricing Step 1 Step 2 Scenario price=Spot price*(U)^2=30*1.1618^2=40.49 Payoff HH=Max(Scenario price-Str